### v1

# Usefulness of Quantum Entanglement for Enhancing Precision in Frequency Estimation

Abstract
We investigate strategies for reaching the ultimate limit on the precision of frequency estimation when the number of probes used in each run of the experiment is fixed. That limit is set by the quantum Cramér-Rao bound (QCRB), which predicts that the use of maximally entangled probes enhances the estimation precision, when compared with the use of independent probes. However, the bound is only achievable if the statistical model used in the estimation remains identifiable throughout the procedure. This in turn sets different limits on the maximal sensing time used in each run of the estimation procedure, when entangled and independent probes are used. When those constraints are taken into account, one can show that, when the total number of probes and the total duration of the estimation process are counted as fixed resources, the use of entangled probes is, in fact, disadvantageous when compared with the use of independent probes. In order to counteract the limitations imposed on the sensing time by the requirement of identifiability of the statistical model, we propose a time-adaptive strategy, in which the sensing time is adequately increased at each step of the estimation process, calculate an attainable error bound for the strategy and discuss how to optimally choose its parameters in order to minimize that bound. We show that the proposed strategy leads to much better scaling of the estimation uncertainty with the total number of probes and the total sensing time than the traditional fixed-sensing-time strategy. We also show that, when the total number of probes and the total sensing time are counted as resources, independent probes and maximally entangled ones have now the same performance, in contrast to the non-adaptive strategy, where the use of independent is more advantageous than the use of maximally entangled ones.
I Introduction
Frequency estimation is of fundamental interest as an essential ingredient for experimental tests of our physical theories [1, 2, 3] and serves as a cornerstone for numerous practical applications, including magnetometers [4, 5], gravimeters [6, 7], and atomic clocks [8, 9, 10]. To estimate an unknown frequency that is encoded in a quantum system, it is necessary to measure an observable of the system and use the resulting data as input for an estimator function. The output of the estimator function represents an estimate of the frequency. The theory of parameter estimation aims to find measurement strategies and estimators that lead to small estimation errors. Quantum Mechanics, on one side, sets limits on how small these erros can be and, on the other side, establishes what are the resources necessary to reach those limits. The ultimate quantum limit on the estimation error is given by the so-called quantum Cramér-Rao bound (QCRB) [11], and the foremost objective in quantum parameter estimation is the development of strategies that can reach this bound.
In this article, we focus our attention on the spectroscopy of two-level systems (qubits), which has an important example in the estimation of the frequency of an atomic transition. The basic metrological scheme to estimating the frequency $\omega$, using $N$ two-level probes, consists in first preparing the $N$ probes in an adequate quantum state, letting them evolve for a time $t$ and then measuring some observable on the final state. The measurement results are fed into an estimator function, which produces an estimate of $\omega$. Specifically, when the $N$ probes are prepared in an initial product state, the QCRB is given by
$\text{Var}_{\omega}(\widehat{\omega}(X_{1},\ldots,X_{N}))\geq\frac{1}{Nt^{2}_{\textrm{prod}}}\,,$ (1)
where $X_{1},\ldots,X_{N}$ is a sample of the outcomes of $N$ independent measurements, each on one of the $N$ probes, $\widehat{\omega}(X_{1},...,X_{N})$ is an unbiased estimator, which satisfies $\text{E}_{\omega}\left[\widehat{\omega}(X_{1},...,X_{N})\right]=\omega$, $t_{\textrm{prod}}$ is the evolution (sensing) time of the product state, and $\rm{Var}_{\omega}\left(\cdot\right)$ is the estimator variance [12, 13, 14]. The inverse of the right hand side in Eq. (1) is equal to the quantum Fisher information, which is a measure of how much information about the parameter $\omega$ can be extracted from the sample $(X_{1},\ldots,X_{N})$. If the results of $\nu$ repetitions of the experiment are taken into account, the QCRB assumes the form
$\text{Var}_{\omega}(\widehat{\omega}(\{X_{1}^{(i)},\ldots,X_{N}^{(i)}\}_{i=1}^{\nu}))\geq\frac{1}{\nu Nt^{2}_{\textrm{prod}}}\,,$ (2)
where ${X_{1}^{(i)},\ldots,X_{N}^{(i)}}$ represents a sample of size $N$, obtained in the experiment $i$. It is well known that, in the asymptotic limit for the number of measurements ($\nu\to\infty$), the maximum likelihood estimator (MLE), derived from a sample of measurements that saturate the quantum information bound (as described below), achieves the saturation of the inequality Eq. (2), thus producing the minimum possible estimation error for this situation[15]. Note that the total time required to satisfy the bound given by Eq. (2) is at least $\nu t_{\textrm{prod}}$, while the total number of probes required is at least $\nu N$. Looking at Eq. (2), it is evident that the error can be reduced by increasing either the number $N$ of probes in the initial product state, the number of times $\nu$ the experiment is repeated, or the evolution (sensing) time $t_{\textrm{prod}}$.
At first sight, since the variance decreases as the inverse of a quadratic polynomial in time and of a linear polynomial in $\nu N$, one can conclude that an efficient way to made the error as small as needed is by increasing $t_{\textrm{prod}}$. Nevertheless, as noted in Ref. [16], there is a limit on how much $t_{\textrm{prod}}$ can be increased because the information on the frequency is generally codified into the relative phase $\phi=(\omega-\omega_{0})t_{\textrm{prod}}$, between the ground state ($\left|g\right\rangle$) and the excited state ($\left|e\right\rangle$) of the probe, where $\omega_{0}$ is a known frequency (the “clock” frequency). Since the relative phase is $2\pi$-periodic, the probabilities of the outcomes of measurements made on the probe become $\pi$-periodic in $\phi$. That is, all phase values $\phi+n\pi$, with $n\in\mathbb{Z}$, produce the same measurement outcomes. This implies that it does not exist a MLE which can make a unique estimation (the likelihood functions are not identifiable) and the Cramér-Rao bound can not be saturated [17, 18]. A solution to this non-identifiability problem is to restrict the value of $\phi$ to $-\pi/2\leq\phi\leq\pi/2$. In other words, since some prior knowledge on the value of the frequency $\omega$ is typically present, if one assumes that this value lies inside some interval around the value of $\omega_{0}$, say $\omega\in(\omega_{0}-\Delta\Omega,\omega_{0}+\Delta\Omega)$, then the sensing time must be restricted to $t_{\textrm{prod}}\leq\pi/(2\Delta\Omega)$. This restriction has the side effect that one cannot choose arbitrary large sensing times to decrease the estimation error.
It is well known that the use of quantum resources may lead to improvement of error bound (2). Preparing the probes in a maximally entangled state, a so-called GHZ state [19], instead of in a product state, gives rise to a smaller QCRB [12, 20]. Specifically, when $N_{\textrm{GHZ}}$ probes are prepared in a initial GHZ state, the quantum Cramér-Rao bound becomes
$\text{Var}_{\omega}\left(\widehat{\omega}(\{X^{(i)}\left(N_{\textrm{GHZ}}\right)\}_{i=1}^{\nu})\right)\geq\frac{1}{\nu N^{2}_{\textrm{GHZ}}t^{2}_{\textrm{GHZ}}}\,,$ (3)
where $t_{\textrm{GHZ}}$ is the sensing time of the GHZ state. It is noteworthy that the variance bound Eq. (3), for GHZ states, scales as $1/N^{2}_{\textrm{GHZ}}$, in contrast to the scaling as $1/N$ of bound (2). This scaling is known as the Heisenberg limit [13] and is the best scaling possible in frequency estimation [12]. When considering multiple probes and equal sensing times, the bound given by Eq.(3) is smaller than the bound in Eq.(2). Therefore, if both bounds can be saturated, using GHZ states offers an advantage over using product states for achieving the smallest possible estimation error. However, the problem of non-identfiability in the likelihood functions plays an important role here. As it will be discussed below, for an initial GHZ state, the information on the frequency $\omega$ will be encoded in the relative phase $\phi_{\textrm{GHZ}}=N_{\textrm{GHZ}}(\omega-\omega_{0})t_{\textrm{GHZ}}$. This phase evolves with the sensing time $t$ faster than the relative phase of an initial product state: $\phi_{\textrm{GHZ}}=N_{\textrm{GHZ}}\cdot\phi$. If, as in the case of an initial product state, one assumes that the value of the frequency $\omega$ lies inside an interval around $\omega_{0}$, say $\omega\in(\omega_{0}-\Delta\Omega,\omega_{0}+\Delta\Omega)$, then , in order to maintain the identifiability of the estimator, the sensing time $t_{\textrm{GHZ}}$ must be restricted to $t_{\textrm{GHZ}}\leq\pi/(2N_{\textrm{GHZ}}\Delta\Omega)$. This means that, for an inicial GHZ state, the maximal sensing time $t_{\textrm{GHZ}}$ must be $N_{\textrm{GHZ}}$ times shorter than the corresponding maximal sensing time $t_{\textrm{prod}}$, for an initial product state. If these restrictions on the sensing times are taken into account in the bounds given in Eqs. (2) and (3), it becomes evident that the advantage of using maximally entangled states, stemming from the quadratic scaling with the number of probes in the estimation precision, can be canceled by the scaling with the inverse of the number of probes of the corresponding maximal sensing time.
Indeed, as will be detailed below, when the estimation procedure consists of a large number of preparation-sensing cycles, with a fixed sensing time, there is no advantage in using maximally entangles states of the probes for improving the precision of frequency estimation, when compared with the use of product states. This result, which may seem counterintuitive, is consequence of the necessity of limiting the sensing time in a way that allows one to get an identifiable statistical model (set of parametrized probability distributions where different values of the frequency must generate different probability distributions) and then a consistent estimator (the estimates converge in probability to the unknown parameter). These conditions are indispensable for saturating the Cramér-Rao bound [21, 22].
The problem of non-identifiability in the statistical model also occurs in interferometry (phase estimation), which is closely related to frequency estimation, and has been discussed by several authors. Ref. [23], in particular, presents a phase estimation procedure with maximally entangles estates of the probes, where the variance of the estimator may scale as $1/N^{2}_{T}$, with $N_{T}$ being the total number of probes used in each run of the experiment. Their procedure, however, requires that, in each run of the experiment, $p$ independent measurements be performed on different sets of $N_{p}$ probes prepared in an initial maximally entangled state, where $N_{p}=1,2,4,...,2^{p-1}$. In the limit of very large $p$, the results of the $p$ measurements can be combined to construct a consistent phase estimator, whose variance scales as $1/N^{2}_{T}$, where $N_{T}=2^{p-1}$ is the total number of probes used in each run of the experiment. The adaptation of this procedure to frequency estimation is straightforward and has been described in Ref. [24] and implemented in Ref. [16]. For phase measurements with the electromagnetic field, Ref. [25] solves the problem of non-identfiability by proposing the use of an adaptive measurement on multiple copies of NOON states distributed in multiple time modes. In their scheme, one first perform $M$ measurements, one on each of $M$ NOON states with the same number $\nu=2^{K}$ of photons. Subsequently, measurements are performed on $M$ NOON states with $\nu=2^{K-1}$, and this sequence of measurements continues on NOON states with $\nu=2^{k}$ for $k=K,K-1,\cdots,1$.
The above schemes successfully get rid of the phase non-identification problem, but their practical implementation is a very challenging task, since a sequence of maximally entangled states with different number of probes has to be prepared and used in each run of the experiment. For this reason, we would like to investigate in this article adaptive estimation strategies that use the same initial state of the probes in each run of the experiment in order to solve the non-identifiability problem in frequency estimation. When $\nu$ experiments are done to estimate a parameter, adaptive methods consist of using the observed data from previous measurements to choose the next quantum measurement, using a suitable cost function. Under the right conditions, such methods surpass the performance of the non-adaptive ones [26, 21]. In fact, adaptive estimation strategies have become a powerful tool for overcoming limitations in diverse estimation problems [21, 27, 28, 29, 30, 22].
Some of the authors have recently demonstrated the benefits of adaptive techniques to overcome the non-identifiability problem for phase estimation. For instance, in the case of single-shot phase estimation in coherent states, an adaptive estimation technique that leverages photon counting, displacement operations, and feedback has been developed. It enables the avoidance of the non-identifiability problem in likelihood functions and, by optimizing the correct cost function, surpasses the performance of Gaussian measurements [30]. Additionally, an adaptive estimation procedure based on confidence intervals has been proposed for the problem of phase estimation in two-level systems. This approach overcomes the non-identifiability limitation of likelihood functions produced by locally optimal measurements and enables the saturation of the QCRB in the asymptotic limit [22].
In this article, we present an adaptive strategy that solves the non-identifiability problem for frequency estimation, in which the probes are prepared in the same initial quantum state in each run of the experiment. This strategy allows one to increase the measurement time in bounds (2) and (3), decreasing the frequency estimation error without non-identification problems. We calculate an error bound for this strategy and discuss how to optimally choose its parameters in order to minimize that bound. Finally, we discuss the possible advantage of using maximally entangled states of the probes, when compared with the use of the probes prepared in an initial product state.
II Identifiability limits: the usefulness of GHZ states
In this section, we show in more detail how the need for an identifiable statistical model cancels any possible advantage of using initial maximally entangled states of the probes for frequency estimation, when the estimation procedure consists of a large number of preparation-sensing cycles with a fixed sensing time.
Any protocol for estimating the transition frequency of a two-level system can be related to a standard Ramsey spectroscopy procedure. First, each one of $N$ two-level probes, of transition frequency $\omega$, is prepared in its ground state $\left|g\right\rangle$ and a Ramsey pulse of frequency $\omega_{0}$ is applied to them. Shape and duration of this pulse are such that each probe is put in a balanced superposition of its ground state $\left|g\right\rangle$ and its excited state $\left|e\right\rangle$. Next, the probes evolve freely for a time $t$ (sensing time) followed by a second Ramsey pulse of same shape and duration of the first pulse. Finally, the internal state of each one is measured and the probability of finding a probe in the excited state $\left|e\right\rangle$ is $P=\cos^{2}[(\omega-\omega_{0})t/2]$. This protocol is repeated $\nu$ times and the measurement results are used to estimate the frequency $\omega$.
In order to be possible to associate each value of $P$ to a single value of $\omega$, the phase $\phi=(\omega-\omega_{0})t/2$ must lie inside an interval of width smaller than half the width of a fringe of $\cos^{2}[(\omega-\omega_{0})t/2]$. Since one typically has some prior knowledge on the value of $\omega$, if it is assumed that this value lies inside some interval around $\omega_{0}$, for example $\omega\in(\omega_{0}-\Delta\Omega,\omega_{0}+\Delta\Omega)$, that condition can be met only if the sensing time $t$ is restricted to $t\leq\pi/(2\Delta\Omega)$. In this case, using the maximum allowed sensing time $t_{\textrm{prod}}=\pi/(2\Delta\Omega)$, the minimum error possible in the estimation of $\omega$ is obtained from the quantum Cramér-Rao inequality (2):
${\rm Var}_{\omega}(\widehat{\omega}(\{X_{1}^{(i)},\ldots,X_{N}^{(i)}\}_{i=1}^{\nu}))\geq\frac{4\Delta\Omega^{2}}{\pi^{2}\nu N}\,,$ (4)
with the use of a total number $\nu N$ of probes. It is natural to consider the total time $T$ of the estimation procedure as a resource. Assuming that the duration of the Ramsey pulses is much shorter than the sensing time $t$ of each run of the experiment, one can set $T=\nu t$. Inserting this relation in Eq. (4), the minimum error in the estimation of $\omega$, when using the maximum sensing time $t_{\rm prob}=\pi/(2\Delta\Omega)$, is determined by
${\rm Var}_{\omega}(\widehat{\omega}(\{X_{1}^{(i)},\ldots,X_{N}^{(i)}\}_{i=1}^{\nu}))\geq\frac{2\Delta\Omega}{\pi TN}\,,$ (5)
showing that the variance of $\omega$ scales as $1/(TN)$. To reach this limit, a total number of at least $\nu N$ probes is needed.
It is well known that, if the probes are put in an initial maximally entangled state, the QCRB for the estimation precision of the frequency $\omega$ becomes smaller, leading, in principle, to a smaller error in the estimation of $\omega$ [12]. For this purpose, each one of $N_{\textrm{G}HZ}$ probes is initialized in its ground state $\left|g\right\rangle$. A Raman pulse is then applied to one of the probes followed by a set of controlled-NOT operations on the other probes. This prepares the probes in a maximally entangled state
$\left|\psi\right\rangle=\frac{1}{\sqrt{2}}\left(\left|gg\cdots g\right\rangle+\left|ee\cdots e\right\rangle\right).$ (6)
Next, the probes evolve freely for a time $t$ followed by a second Ramsey pulse on the same probe that received the fist pulse and a set of controlled-NOT operations on the other probes. Finally, the internal state of the probe that received the Ramsey pulses is measured and the probability of finding it in the excited state $\left|e\right\rangle$ is $P=\cos^{2}[N_{\textrm{G}HZ}(\omega-\omega_{0})t/2]$. This protocol is repeated $\nu$ times and the measurement results are used to estimate the frequency $\omega$ [13].
Notice that the probability $P$ oscillates $N_{\textrm{G}HZ}$ times faster than the probability corresponding to initial product states, leading to the improved quantum Cramér- Rao bound (3). Comparing bounds (2) and (3), one can see that, for the same sensing time $t$ and equal numbers $\nu$ of runs, the use of $N$ probes in an initial maximally entangled state may lead to an improvement in the precision of the estimation of $\omega$ by a factor $\sqrt{N}$. This clearly shows the advantage of the use of entangled states for improving the bound on the estimation of the frequency $\omega$. However, in order to maintain the phase $\phi_{\textrm{G}HZ}=N_{\textrm{G}HZ}(\omega-\omega_{0})t/2$ inside an interval of width smaller than a half width of a fringe of $P$, the sensing time $t$ must be limited. If one assumes, like in the case of an initial product state of the probes, that the frequency $\omega$ lies inside the interval $(\omega_{0}-\Delta\Omega,\omega_{0}+\Delta\Omega)$, the sensing time must be restricted to $t\leq\pi/(2N_{\textrm{G}HZ}\Delta\Omega)$. Using the maximum allowed sensing time $t_{\textrm{G}HZ}=\pi/(2N_{\textrm{G}HZ}\Delta\Omega)$ in bound (3), results in
${\rm Var}_{\omega}\left(\widehat{\omega}(\{X^{(i)}\left(N_{\textrm{GHZ}}\right)\}_{i=1}^{\nu})\right)\geq\frac{4\Delta\Omega^{2}}{\pi^{2}\nu}\,.$ (7)
This bound does not even depend on the number $N_{\textrm{G}HZ}$ of probes in the entangled initial state. This is a consequence of the fact that, to guarantee a consistent estimator, it is not possible to increase the number $N_{\textrm{GHZ}}$ of probes in the entangled state without decreasing the sensing time $t_{\textrm{G}HZ}$ by the same proportion.
Comparison of bounds (7) and (4) shows that the use of initial maximally entangled states is disadvantageous when compared with the use initial product states of the probes, if the only resource taken into account is the total number $\nu N=\nu N_{\textrm{G}HZ}$ of probes. Indeed, the maximal reachable precision in the estimation of $\omega$ is worse by a factor $\sqrt{N}$
Since the sensing time $t$, when using probes in a product state, can be much larger than the sensing time when using maximally entangled states, it is essential to count the total time $T=\nu t$ of the experiment as a resource. In this case, bound (7) becomes
${\rm Var}_{\omega}\left(\widehat{\omega}(\{X^{(i)}\left(N_{\textrm{GHZ}}\right)\}_{i=1}^{\nu})\right)\geq\frac{2\Delta\Omega}{\pi TN_{\textrm{GHZ}}}\,.$ (8)
Notice that the total number of probes necessary to reach this limit is at least $\nu N_{\textrm{GHZ}}$. Bound (8) should be compared with bound (5) for initial product states. For equal total duration $T$ of the estimation procedures and equal number $N=N_{\textrm{G}HZ}$ of probes used in each run, they set the same upper limit to the precision of the estimation of $\omega$. However, in each run of the estimation process, the maximum sensing time $t_{\textrm{G}HZ}$ allowed with the use of $N_{\textrm{G}HZ}$ entangled probes is shorter than the maximum sensing time $t_{\textrm{p}rob}$, optimal for the use of $N=N_{\textrm{G}HZ}$ independent probes, by a factor $N_{\textrm{G}HZ}$. Since $T$ is the same in both cases, this implies that the number $\nu$ of runs with the use of entangled probes has to be $N_{\textrm{G}HZ}$ times larger than the number $\nu$ of runs with the use of independent probes. This, on the other side, implies that the total number $\nu N$ of probes used in the former case is larger than the total number of probes used in the latter case. In the limit of large number of measurements, $\nu\gg 1$, the two bounds are saturated and set the best precision effectively reachable in both cases. Consequently, when the total number $\nu N$ of probes and the total duration $T$ of the estimation process are counted as fixed resources, the use of initial GHZ states of the probe is disadvantageous when compared with the use of independent probes. In order to reach the same precision within a fixed total time $T$, the strategy that uses GHZ states needs $N_{\textrm{G}HZ}$ more probes than the strategy that uses product states.
The above discussion makes it clear that, if the initial state of the probes and the sensing time are fixed, the need for an identifiable statistical model cancels any advantage of the use of entangled states for estimation of the frequency $\omega$. In order to address the limitation on the sensing time, we shall introduce an adaptive estimation strategy that takes advantage of the fact that the maximum sensing time increases as $1/\Delta\Omega$ when the frequency interval shrinks. The key idea of this approach is to start with a small number of identical preparation-sensing-measurement cycles to produce an initial estimate of $\omega$. This first estimation allows one to shorten the interval $\Delta\Omega$ and, consequently, to increase the sensing time for the next set of measurements. This process is repeated until the desired estimation error is achieved (see Fig. 1). We present the details of this strategy in the next sections.
Figure 1: Adaptive-time frequency estimation. The first step consist of a measurement over a phase codified in a quantum state. If the error in the measurement is smaller than the step goal, the measurement time is incremented, the goal is updated and the process is repeated, in this case, the quantum information for the next measurment is larger. If the error in the measurement is equal or larger than the step goal, the measurement is repeated without incrementing the measurement time. This procedure is repeated until the global error goal is achieved.
III Estimation with Confidence Intervals
In this section, we review frequency optimal measurements and the conditions that lead to consistent estimators to propose and simulate a frequency estimation strategy that minimizes the error by increasing the measurement time. We call the proposed method Adaptive-Time Frequency Estimation (ATFE).
III.1 Optimal measurements and asymptotic consistent estimators
The goal is to estimate an unknown parameter $\omega$, codified into a two-level system as a phase $\phi=(\omega-\omega_{0})t$, where $t$ and $\omega_{0}$ are control parameters. Without loss of generality, the codification process can be represented by an unitary operation on a fiducial state of the two-level system
$\rho(\omega,t)=\hat{U}_{\omega}(t)\rho\,\hat{U}^{\dagger}_{\omega}(t),$ (9)
with $\hat{U}_{\omega}(t)=e^{-i(\omega-\omega_{0})t\frac{\vec{\sigma}}{2}\cdot\vec{n}}$. Here, $\vec{\sigma}$ is the vector of Pauli matrices, $\vec{n}$ a unit vector and $\rho=\frac{1}{2}\left(I+\vec{\sigma}\cdot\vec{a}\right)$ a fiducial state of the two-level system, characterised by the Bloch vector $\vec{a}$. The codification process leads to a new state, which is represented by the Bloch vector resulting from the rotation of $\vec{a}$, by the angle $(\omega-\omega_{0})t$, around an axis parallel to $\vec{n}$
To estimate $\omega$ we use an estimator $\widehat{\omega}:\mathcal{X}\to\Omega$, which is a function whose domain is defined in the set of measurement outcomes $\mathcal{X}$. Since the frequency is codified in the relative phase between the two levels of the system, $\rho(\omega,t)$ is periodic in $\omega$, and as a result, the estimations would also be periodic in $\omega$. To ensure uniqueness in the estimation, the range of the estimator’s values should be defined inside an interval smaller than that period.
The general description of measurements on a quantum system is given by Positive Operator Valued Measures (POVMs) [31, 32]. Given an outcome space $\mathcal{X}$ and the Borel $\sigma$-algebra $\mathcal{B}\left(\mathcal{X}\right)$ that represents the events that can be observed in an experiment, a POVM with outcome space $\mathcal{X}$ is a $\sigma$-additive map (countable additivity for a sequence of pairwise disjoint events) $P:\mathcal{B}\left(\mathcal{X}\right)\to\mathcal{B}\left(\mathcal{H}\right)$ from the Borel $\sigma$-algebra to the space $\mathcal{B}\left(\mathcal{H}\right)$ of bounded operators on $\mathcal{H}$ [31, 33]. In the case of a finite outcome space $\mathcal{X}$, the set of positive operators $P(k)$ on $\mathcal{H}$ with the property that $\sum_{\mathcal{X}}P(k)=I$, where $I$ is the identity operator on $\mathcal{H}$, determines a positive operator valued measure (POVM). Hence, given a measurement on a state $\rho(\omega)$ of a quantum system, with outcomes $x\in\mathcal{X}\subset\mathbb{R}$ and described by a POVM $P=\left\{P(x)\mid x\in\mathcal{X}\right\}$, the conditional probability distribution for $x$ is given by Born’s rule
$p(x\mid\omega)=\textrm{Tr}\left[P(x)\rho(\omega)\right].$ (10)
Thus, given a sample $X$ of results from the application of the POVM $P$, the expected value of any estimator $\widehat{\omega}\left(X\right)$ based on this sample is defined as
$\text{E}_{\omega}\left[\widehat{\omega}\left(X\right)\right]=\sum_{x\in\mathcal{X}}p(x\mid\omega)\widehat{\omega}(x).$ (11)
It can be the case that the estimator is a multivalued function. For instance, as noted above, the frequency estimator that uses measurements from a two-level system is periodic. This periodicity introduces a problem when calculating the estimation error. For example, given $t$, if $\hat{\omega}_{P}$ is the period in frequency of the quantum state (9), two possible estimations for the same experimental data are $\omega+\epsilon$ and $\omega+\epsilon+m\,\hat{\omega}_{P}$, with $m$ an integer and $\epsilon\ll 1$; one of these estimations has an error $\epsilon$, whereas the other has an error $\epsilon+m\,\hat{\omega}_{P}$.
To correctly calculate that variance, it is necessary to modify the cost function to account for the fact that estimations that differ by a multiple of the period correspond to the same value of the parameter to be estimated. In such scenarios, an appropriate measure is given by the Holevo variance of $\widehat{\omega}(X)$ [31, 29]:
$\textrm{Var}_{\omega}\left[\widehat{\omega}(X)\right]=(\frac{\hat{\omega}_{P}}{2\pi})^{2}\Big{(}\Big{\lvert}\sum_{x\in\mathcal{X}}p(x\mid\omega)e^{2\pi\mathrm{i}(\widehat{\omega}(x))/\hat{\omega}_{P}}\Big{\rvert}^{-2}-1\Big{)}\,.$ (12)
This variance reduces to the usual definition when the error is small.
When a POVM $P=\left\{P(x)\mid x\in\mathcal{X}\right\}$ is performed on the state $\rho(\omega)$, the minimum attainable estimation error of any unbiased periodic estimator $\widehat{\omega}\left(X\right)$ of the frequency $\omega$ is bounded by the Cramér-Rao bound (CRB)
$\textrm{Var}_{\omega}\left[\widehat{\omega}\left(X\right)\right]\geq\frac{1}{F(\omega;P)}.$ (13)
Here, $F(\omega;P)$ represents the Fisher information that a sample $X$ of results from the POVM $P$ carries about the parameter $\omega$. It is computed from the probability distribution of $X$ [31, 18]:
$F(\omega;P)=\textrm{E}_{\omega}\left[\left(\frac{\partial}{\partial\omega}\log\left(p(x\mid\omega)\right)\right)^{2}\right].$ (14)
Particularly, for a sample $X_{1},\ldots X_{\nu}$ of size $\nu\geq 1$, obtained from $\nu$ identically and independently applied measurements $P$, the CRB for any unbiased estimator $\widehat{\omega}_{\nu}:=\widehat{\omega}\left(X_{1},\ldots,X_{\nu}\right)$ based on this sample is given by
$\textrm{Var}_{\omega}\left[\widehat{\omega}_{\nu}\right]\geq\frac{1}{\nu F(\omega;P)}\,.$ (15)
This inequality follows from the additivity of the Fisher information [34]. The estimators that saturate the CRB are called efficient, and when this condition is met in the asymptotic limit ($\nu\to\infty$), they are called asymptotically efficient. A well-known result in statistics is that, under a set of regularity conditions, the MLE is an asymptotically efficient unbiased estimator [35, 21].
The ultimate limit of precision in quantum mechanics is achieved by optimizing the quantity $F(\omega;P)$ over all POVMs. This optimization process yields the quantum Fisher information (QFI) about the unknown parameter $\omega\in\Omega\subseteq\mathbb{R}$. The QFI is a function of $\omega$ that is independent of any specific POVM and is defined as [31, 32]:
$F_{\textrm{Q}}(\omega)=\text{Tr}\left[\rho(\omega)\lambda(\omega)^{2}\right]\,,$ (16)
where $\lambda(\omega)$ is the symmetric logarithmic derivative (SLD), which is implicitly defined by the equation
$\frac{d\rho(\omega)}{d\omega}=\frac{1}{2}\left(\lambda(\omega)\rho(\omega)+\rho(\omega)\lambda(\omega)\right)\,.$ (17)
The QFI provides a generalization of the Cramér-Rao bound to the quantum domain, the so-called quantum Cramér-Rao bound (QCRB)
$\textrm{Var}_{\omega}\left[\widehat{\omega}(X)\right]\geq\frac{1}{F_{\rm{Q}}(\omega)}\,.$ (18)
Additionally, the Fisher information of a POVM $P$ satisfies [11]
$F(\omega;P)\leq F_{\textrm{Q}}(\omega)\,.$ (19)
This inequality is known as the quantum information bound (QIB). The POVMs that saturates Eq. (19) are called optimal and are the most sensitive measurements for the estimation of the parameter. A sufficient condition for achieving the QIB is given by the POVM $P_{\rm{L}}(\omega)$, whose elements are the projectors onto the eigenspaces of the SLD operator [11, 15]. By construction, this POVM depends on $\omega$, which is the parameter to be estimated (it is locally optimal). For this reason, if one considers $\left\{P_{\rm{L}}(g)\right\}_{g\in\Omega}$ as a one parameter family of POVMs, one can only guarantee that $P_{\rm{L}}(g)$ achieves the QIB if $g=\omega$, turning the method useless because the value of $\omega$ is unknown. However, in some systems, one can find initial conditions where any POVM $P_{\rm{L}}(g)$ saturates Eq. (19) for every $g\in\Omega$ and independently of $\omega$ [15].
In the context of frequency estimation, the quantum Fisher information (QFI) of the state Eq. (9) is independent of $\omega$ and is given as:
$F_{\rm{Q}}=t^{2}\left[1-\left(\vec{a}\cdot\vec{n}\right)^{2}\right]\,.$ (20)
When the Bloch vector associated with the initial state of the probe is perpendicular to the rotation axis of the frequency codification process, $\vec{n}\cdot\vec{a}=0$, the quantum Fisher information $F_{\rm{Q}}$ reaches its maximum value $t^{2}$. Moreover, under this initial conditions, any POVM $P_{\rm{L}}(g)$ saturates Eq. (19) for any $\omega\in\Omega$ [36, 22]. We will assume this initial condition from now on. Explicitly, the elements for any $P_{\rm{L}}(g)=\left\{P(x;g)\mid x\in\left\{0,1\right\}\right\}$ are
$\begin{split}&P(0;g)=\frac{1}{2}\left(I+\vec{n}\times\vec{a}(g)\cdot\vec{\sigma}\right),\\ &P(1;g)=\frac{1}{2}\left(I-\vec{n}\times\vec{a}(g)\cdot\vec{\sigma}\right),\end{split}$ (21)
where $a(g)=\cos((g-\omega_{0})\,t)\vec{a}+\sin((g-\omega_{0})\,t)\vec{n}\times\vec{a}$, $g\in\Omega$. Thus, according to the Born’s rule
$p(x\mid\omega;t)=\begin{cases}\frac{1}{2}\left[1+\sin((\omega-g)\,t)\right]\,\text{if }x=0\\ \frac{1}{2}\left[1-\sin((\omega-g)\,t)\right]\,\text{if }x=1\end{cases}$ (22)
and from Eq. (14),
$F(\omega;P_{\textrm{L}}(g))=t^{2}=F_{\textrm{Q}}(\omega),\,\,\,\forall g,\omega\in\Omega.$ (23)
To exemplify how a physical measurement is codified in the POVM (21) we use a specific example considering that a two-level system is algebraically equivalent to the electron spin. Let the initial condition of the probe be a state with spin in the $x$ direction, described by the Bloch vector $a=(1,0,0)$, and the codification process be a rotation around the $y$ direction, $\vec{n}=(0,1,0)$. If $g=\omega_{0}$, from Eq.  (21) it is easy to see that the elements of the POVM $P_{\rm{L}}(g)$ describe a measurement of the component of the spin in the $z$ direction.
As noted above, the MLE obtained by a sequence of $\nu$ independent outcomes of any $P_{\rm{L}}(g)$ can saturate the CRB under a set of regularity condition [34]. For frequency estimation, one of the regularity conditions that fails to be satisfied is that the likelihood function has to be identifiable. For example, using the POVM (21) the likelihood function is
$L(\omega)=p(0\mid\omega;t)^{m}\left(1-p(0\mid\omega;t)\right)^{\nu-m},$ (24)
where $m$ is the number of 00s in the measured data. Inserting Eq. (22) in the above expression, it is straightforward to see that this likelihood function has multiple global maxima, separated of each other by $\pi$, rendering it asymptotically inconsistent and biased. A way to achieve saturation of the CRB is to render the likelihood function (24) identifiable by restricting its image to a sufficiently small interval where there is only one global maximum.
Summarizing, the POVM (21) saturates the QIB, Eq. (19), but two problems arise if one wants to use it to minimize the estimation error: i) it is locally optimal and ii) for large measurement times its associated likelihood, Eq. (24), becomes non-identifiable. The first problem can be avoided using the optimal initial conditions or adaptive estimation schemes [21]; the second one can be avoided by requiring an unique estimation in $\omega\in\Omega$, which can be accomplished if $t\leq\pi/(2\Delta\Omega)$, where $2\Delta\Omega$ is the length of the interval defined by $\Omega$. Fig. 2 shows an example of how the number of maxima increases by one if $t$ increases by $\pi/(2\Delta\Omega)$, making it impossible to find a consistent frequency MLE for $t>\pi/(2\Delta\Omega)$.
Non-identifiability problems arises in qubit phase estimation, where the locally optimal POVM produces a likelihood with two maxima. In [22] this non-identifiability was solved using an adaptive estimation technique based on the construction of confidence intervals from a prior sample of the canonical phase measurement [37, 38] and a posterior adaptive sequence implementation of the locally optimal POVM. One of the objectives of this paper is to find an adaptive estimation strategy that allows one to increase the measurement time in the bound (2), decreasing the frequency estimation error in two-level systems.
To introduce the proposed strategy, we first define confidence intervals as a set of plausible values that are likely to contain the true parameter value, with the confidence level representing the proportion of such intervals containing the parameter’s value, in the limit of an infinite number of repeated experiments. Using the MLE $\widehat{\omega}_{\textrm{MLE}}(X)$ for a sample $X$, an estimator for a confidence interval with confidence level $0\leq C_{l}=1-\alpha\leq 1$ can be computed as follows:
$\widehat{\textrm{CI}}(\widehat{\omega}_{\textrm{MLE}}(X))=\left(\widehat{\omega}_{\textrm{MLE}}(X)-z_{\alpha/2}\cdot F(\widehat{\omega}_{\textrm{MLE}}(X))^{-\frac{1}{2}},\widehat{\omega}_{\textrm{MLE}}(X)+z_{\alpha/2}\cdot F(\widehat{\omega}_{\textrm{MLE}}(X))^{-\frac{1}{2}}\right)\ ,$ (25)
where $F(\omega)$ is the Fisher information of the sample, $z_{\alpha}$ is the $\alpha$-ith quantile of the standard normal distribution ($P(Z\geq z_{\alpha})=\alpha$) [35]. Hence, every outcome of $\widehat{\textrm{CI}}(\widehat{\omega}_{\textrm{MLE}}(X))$ is a possible confidence interval.
Eq. (25) assumes that the distribution of the MLE is asymptotically normal, with its mean at $\omega$ and variance equal to the inverse of the Fisher information. This requires the use of estimation strategies with consistent estimators. Thus, the identifiability of the parameters becomes crucial, since it is both a sufficient and necessary condition for obtaining an asymptotically consistent MLE [35, 17].
Figure 2: Normalized log-likelihood functions produced by $P_{\rm{L}}(\tilde{g})$. When $\tilde{t}=1/4$ the likelihood function is identifiable. For $\tilde{t}=1/2$ and $\tilde{t}=3/4$ there are $2$ and $3$ maxima, respectively, making the likelihood functions not identifiable. For all cases, the likelihood functions were plotted using $64$ data from $P_{\rm{L}}(\tilde{g})$ and assuming that $\tilde{\omega}=2$.
III.2 Minimizing the estimation error using adaptive measurements and confidence intervals: Adaptive-Time Frequency Estimation (ATFE)
A quick explanation of the method is as follows: we start by finding a set $\Omega$ such that $\omega\in\Omega$ and setting a measurement time $t_{1}$ such that the likelihood function given in Eq. (24) has a single stationary point as a global maximum within that region. Subsequently, we perform an optimal POVM to obtain the data $x_{1}$ and an estimation $\widehat{\omega}_{\textrm{MLE}}(x_{1})$. By using this estimate, a confidence interval $CI_{1}=\widehat{\textrm{CI}}(\widehat{\omega}_{\textrm{MLE}})(x_{1})\subset\Omega$ is obtained, according to Eq. (25). Assuming that $\omega\in CI_{1}$, we set a new measurement time $t_{2}>t_{1}$ such that the statistical model restricted to $CI_{1}$ remains identifiable. After repeating the experiment, a new frequency estimation inside a smaller confidence interval $CI_{2}$ is obtained. By iteratively performing this procedure, we increase the measurement time at each adaptive step until the desired error is achieved.
To describe the method in detail, one first assumes that $\Omega=[\omega_{0}-\Delta\Omega,\omega_{0}+\Delta\Omega]$ and defines an adimensional time, $\tilde{t}=t\Delta\Omega/2\pi$, with the goal to estimate the adimensional frequency $\tilde{\omega}=(\omega-\omega_{0})/\Delta\Omega$. With this notation $\tilde{\omega}\in\tilde{\Omega}=[-1,1]$. For $N$ probes initially prepared in a product state, the choice of the maximal measurement time $\tilde{t}=1/4$ guarantees that
$-\pi/2\leq(\omega-\omega_{0})t\leq\pi/2\,,$ (26)
whereas for $N$ probes initially prepared in a GHZ state, the above condition is satisfied when $\tilde{t}=1/(4N)$. The relation between the variance of a estimator for the adimensional frequency $\tilde{\omega}$ and an estimator for the actual frequency $\omega$ is
$\textrm{Var}\left(\widehat{\omega}\right)=\Delta\Omega^{2}\cdot\textrm{Var}\left(\widehat{\tilde{\omega}}\right)\,.$ (27)
For simplicity, throughout the rest of the paper, $\widehat{\omega}$ will be an estimator for $\tilde{\omega}$.
For a single probe, one starts by selecting a measurement time $\tilde{t}_{1}=1/4$. This yields a likelihood function with a single global maximum in $\tilde{\Omega}$, when the parameter $g$ in the POVM Eq. (21) is close to $\tilde{\omega}$. Then, the first estimation strategy, termed $M_{1}$ and described in detail in the following, is employed. This strategy consists of making measurements on $\nu_{1}$ independent probes, each of them producing a Fisher information $F_{Q_{1}}=(2\pi\tilde{t}_{1})^{2}=\pi^{2}/4$, until the total fisher information $F_{1}(\tilde{\omega})$, obtained after the $\nu_{1}$ measurements, satisfies $z_{\alpha/2}\cdot F(\widehat{\omega}_{\textrm{MLE}}(x_{1}))^{-\frac{1}{2}}\approx 1/2$, where $x_{1}$ is the data set containing the measurements results. The confidence interval for the outcome $x_{1}$ is then (see (25)) $CI_{1}=\left(\widehat{\omega}_{\textrm{MLE}}(x_{1})-1/2,\widehat{\omega}_{\textrm{MLE}}(x_{1})+1/2\right)$, where $\widehat{\omega}_{\textrm{MLE}}(x_{1})$ is the estimation using the strategy $M_{1}$. The subsequent step is to extend the measurement time to $\tilde{t}_{2}=2\tilde{t}_{1}$, thereby increasing the Fisher information that can be gained in each measurement. Notice that increasing the measurement time introduces a second maximum in $\tilde{\Omega}$ that is displaced by $1$ with respect to the first maximum. This second maximum can turn the likelihood function non identifiable inside the set $\tilde{\Omega}$. However, if the first strategy was successful, this second maximum lies outside $CI_{1}$. Therefore, restricting the next estimation strategy $M_{2}$ to the set $CI_{1}$ guarantees that the likelihood function remains identifiable. The probability that the parameter is not in $CI_{1}$ is $1-P(\theta\in CI_{1})=\alpha$. In general, the $i$th strategy $M_{i}$ consists of $\nu_{i}$ measurements that generates a sample $X_{i}$, and $\widehat{\omega}_{\textrm{MLE}}(X_{1},X_{2},\ldots,X_{i})$. By choosing the Fisher information of the $M_{i}$ strategy such that $z_{\alpha/2}\cdot F(\widehat{\omega}_{\textrm{MLE}}(X_{1},X_{2},\ldots,X_{i}))^{-\frac{1}{2}}\approx 1/(i+1)$ we obtain an estimator for the confidence interval
$\begin{split}&\widehat{CI}_{i}=\\ &\left(\widehat{\omega}_{\textrm{MLE}}(X_{1},X_{2},\ldots,X_{i})-1/(i+1),\right.\\ &\widehat{\omega}_{\textrm{MLE}}(X_{1},X_{2},\ldots,X_{i})+1/(i+1)\left.\right)\,,\end{split}$ (28)
which allows one to increase the measurement time to $\tilde{t}_{i}=i\tilde{t}_{1}$ for the next measurement strategy, increasing the Fisher information gained in each measurement to $F_{Q_{i}}(\tilde{\omega})=(\pi i)^{2}/4$ and diminishing the estimation error. In order to find the bound for the estimation error of the whole procedure we have to specify the estimation strategies $M_{i}$.
To obtain a performance close to the QCRB, it is necessary to obtain an asymptotic efficient estimator (asymptotically unbiased, with variance equal to the inverse of Fisher information). To this end, we use the Adaptive Quantum State Estimation method (AQSE) [21], in the region $CI_{i}$, with the family of POVMs $\left\{P_{\textrm{L}}(\tilde{g})\right\}_{\tilde{g}\in\tilde{\Omega}}$, as the $i$-th optimally local estimation strategy $M_{i}$ [39, 21], with $\tilde{g}=(g-\omega_{0})/\Delta\Omega$. Notice that each strategy $M_{i}$ comprises $\nu_{i}$ measurement steps. The AQSE begins with an arbitrary initial guess $\tilde{g}_{0}$. At this point, the locally optimal measurement $P_{\rm{L}}(\tilde{g}_{0})$ is applied. Assuming that the outcome $x_{1}$ is observed, one obtains the likelihood function $L_{1}(\tilde{\omega};x_{1};\tilde{g}_{0})=p(x_{1}\mid\tilde{\omega};\tilde{g}_{0})$ and applies the MLE to produce an estimate $\tilde{g}_{1}=\widehat{\omega}_{1}(x_{1})=\operatorname*{arg\,max}_{\tilde{\omega}\in\Omega}L_{1}(\tilde{\omega};x_{1};\tilde{g}_{0})$, which will be used as the new guess for the POVM $P_{\rm{L}}$. Thereby, for the step $n\geq 2$, one applies the POVM $P_{\textrm{L}}(\tilde{g}_{n-1}=\widehat{\omega}_{n-1}(x_{1},...,x_{n-1}))$, where $\widehat{\omega}_{n-1}(x_{1},...,x_{n-1}))$, is the estimation from the previous stage, which used the outcomes $x_{1},...,x_{n-1}$. If the outcome $x_{n}$, is observed, the likelihood function at $x_{n}$ for step $n$ is
$L_{n}(\tilde{\omega};x_{1},...,x_{n};\tilde{g}_{n-1})=\prod_{i=1}^{n}p(x_{i}\mid\tilde{\omega};\tilde{g}_{i-1})\,,$ (29)
from which one gets the nth guess $\tilde{g}_{n}=\widehat{\omega}_{n}\left(x_{1},...,x_{n}\right)$ by applying the MLE:
$\widehat{\omega}_{n}(x_{1},...,x_{n})=\operatorname*{arg\,max}_{\tilde{\omega}\in\tilde{\Omega}}L_{n}(\tilde{\omega};x_{1},...,x_{n};\tilde{g}_{n-1})\,.$
Since $L_{n}$ is close to zero when $n\gg 1$, it is more convenient to use the natural logarithm of the likelihood function, which has the same maxima than the likelihood function. We denote the logarithm of the likelihood function by $l(\tilde{\omega})$. An equivalent definition for the MLE of $\tilde{\omega}$, given a sample $(X_{1},...,X_{n})$, is given by
$\widehat{\omega}_{n}(X_{1},...X_{n})=\operatorname*{arg\,max}_{\tilde{\omega}\in\tilde{\Omega}}l(\tilde{\omega};X_{1},...,X_{n}).$ (30)
Since, for the optimal initial condition ($\vec{n}\cdot\vec{a}=0$), the Fisher information $F_{Q_{i}}$ of each measurement step of the $i$th strategy does not depend on the parameter $\tilde{g}$, the CRB for $\widehat{\omega}_{n}(X_{1},...X_{n})$, produced from the AQSE method in $n$ adaptive steps, is
$\text{Var}_{\omega}\left[\widehat{\omega}_{n}(X_{1},...,X_{n})\right]\geq\frac{1}{nF_{Q_{i}}}=\frac{1}{n(2\pi\tilde{t}_{i})^{2}},$ (31)
for each strategy $M_{i}$.
It could be argued that AQSE is not necessary because the Fisher information is the same regardless of the value of $\tilde{g}$. However, AQSE is needed because it ensures a asymptotically normal estimator [21, 22], which is a necessary condition to justify the use of Eq. (25), which define our confidence intervals. Adopting AQSE as the $M_{i}$ estimation strategy implies that the protocol of estimation with confidence intervals consists of two adaptive processes: the primary adaptive process, with total Fisher information $F_{i}$, which increases the measurement time after a certain confidence interval is attained, and the secondary adaptive process (with a fixed measurement time $\tilde{t}_{i}$ and Fisher information $F_{Q_{i}}=(2\pi\tilde{t}_{i})^{2}$ for each measurement), which involves the implementation of AQSE as the $M_{i}$ estimation strategy, wherein the POVM is altered after each measurement until the required confidence interval is achieved.
We determine now the estimation error of ATFE. For a given confidence level $C_{l}\in\left[0,1\right]$, in each adaptive step, we have two contributions to the estimation error: i) when the confidence interval $CI=(\widehat{\omega}-E,\widehat{\omega}+E)$, where $E$ is the marginal error and $\widehat{\omega}$ is the estimation, contains the value of the parameter and ii) when it does not. When the confidence interval contains the value of the parameter, the MLE from AQSE converges in distribution to a normal distribution with mean at $\tilde{\omega}$ and variance equal to the inverse of the Fisher information [35, 21]. On the other hand, if the confidence interval does not contain the parameter, AQSE produces an error larger than $E$ [22].
Thereby, given that $(C_{l})^{i}$ and $(1-C_{l})\sum_{j=1}^{i}(C_{l})^{j-1}$ are the probability that the parameter is inside and outside the confidence interval at primary adaptive step $i$, respectively, the estimation error for a total number $\nu=\sum_{i}^{S}\nu_{i}$ of measurements is
$\begin{split}&\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]\geq\frac{(C_{l})^{S}}{\sum_{i=1}^{S}\nu_{i}F_{Q_{i}}}\\ &+\left(1-C_{l}\right)\left(E_{1}^{2}+C_{l}E^{2}_{2}+\cdots+(C_{l})^{S-1}E^{2}_{S}\right),\end{split}$ (32)
where $S$ is the total number of primary time-adaptive steps, and $\nu_{i}$, the number of measurements using time $\tilde{t}_{i}$ in the strategy $M_{i}$, is set by the requirement that the resulting confidence interval be $CI_{i}$. When $\nu\gg 1$, the second contribution dominates in Eq. (32), since performing additional measurements does not reduces the error if the parameter is outside of the confidence interval. For this protocol to be useful, the confidence levels should be chosen in such a way that the second term of Eq. (32) is negligible. The marginal error is a function of the Fisher information and the quantile of the standard normal distribution. To estimate the minimum number $\nu_{i}^{\textrm{min}}$ of measurements at step $i$, required to produce a confidence interval of length less than the marginal error $E_{i}$, we use Eq. (25). From that equation it follows that $E^{2}_{i}\geq z^{2}_{\alpha/2}/\sum_{j=1}^{i}\nu_{j}F_{Q_{j}}$, where $F_{Q_{j}}$ is the Fisher information of each AQSE measurement of the estimation strategy $M_{j}$. Since $F_{Q_{j}}=\pi^{2}j^{2}/4$ and $E_{i}=1/(i+1)$, this expression leads to the condition $\sum_{j=1}^{i}\nu_{j}j^{2}\geq\frac{4}{\pi^{2}}z^{2}_{\alpha/2}\left(i+1\right)^{2}$, which, for $i\geq 2$, can be rewritten as $\nu_{i}^{\textrm{min}}i^{2}\geq\frac{4}{\pi^{2}}z^{2}_{\alpha/2}\left(i+1\right)^{2}-\sum_{j=1}^{i-1}\nu_{j}^{\textrm{min}}j^{2}$. If one substitutes $\sum_{j=1}^{i-1}\nu_{j}^{\textrm{min}}j^{2}$, on the right-hand side of the former expression, by $\frac{4}{\pi^{2}}z^{2}_{\alpha/2}i^{2}$, one finally gets
$\nu_{i}^{\textrm{min}}\geq\frac{4}{\pi^{2}}z^{2}_{\alpha/2}\frac{2i+1}{i^{2}}.$ (33)
Figure 3: Lower bound in the Holevo variance for the maximum likelihood estimator produced by ATFE. Eq. (32) is plotted for the confidence levels $C_{l}=0.95$ (yellow line), $C_{l}=0.99$ (blue line), and $C_{l}=0.999$ (green line). The cross markers in the curves indicate the adaptive step when the time is increased in the method.
The right-hand side of Eq. (33) diminishes as a function of $i$ and will eventually become smaller than $1$ as the number of primary time-adaptive steps increases. If we call $S_{1}$ the number of steps for which this happens, then, from Eq. (33), it is easy to see that $S_{1}\approx\frac{8}{\pi^{2}}z^{2}_{\alpha/2}$. As examples of typical values of $S_{1}$, for a confidence level $C_{l}=0.99$, the value of $S_{1}$ will be $S_{1}\approx 5$, whereas for a confidence level $C_{l}=0.999999$, $S_{1}$ will increase to $S_{1}\approx 19$. If one additionally assumes that the total number of measurements at step $S_{1}$ is such that the MLE is already a normal estimator that saturates the Crámer-Rao bound, then, when $i>S_{1}$, the choice of $\nu_{i}=1$ will produce an error smaller than the marginal error goal for that step. This means that the measurement time can be increased by $\tilde{t}_{1}$ at each measurement, after step $S_{1}$. The behavior of the lower bound Eq. (32) multiplied by the total number $\nu$ of measurements is shown in Fig. 3 for increasing confidence levels of the confidence intervals. For comparison, the solid blue line shows the QCRB for a fixed measurement time $\tilde{t}_{i}=1/4$, which corresponds to the minimum possible error if no adaptive changes in the measurement times are allowed ( cf. Eq. (2)). In this case $\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]=4/(\pi^{2}\nu)$. From that figure it can be seen that, as long as the second term of Eq. (32) is negligible, the ATFE performs much better than the non-adaptive method. That term, which is almost constant and limits the minimum error reachable in the ATFE, depends only on the confidence level $C_{l}$ and on the confidence interval widths $E_{i}$. Its value decreases when the confidence level $C_{l}$ increases; this, in turn, increases the values of $\nu$ for which its contribution becomes important. In section III.3, we simulate the measurement and show that the protocol reaches this bound.
Now we focus on obtaining analytical approximations to Eq. (32) to better understand how the error diminish as a function of the resources. First, notice that the total number of measurements at step $S$, $\nu=\sum_{i=1}^{S}\nu_{i}$, can be approximated by $\nu=\sum_{i=1}^{S}\nu_{i}^{\textrm{min}}$. Using Eq. (33) for $i\leq S_{1}$ and $\nu_{i}^{\textrm{min}}=1$ for $i>S_{1}$, one obtains
$\nu=S_{1}\sum_{i=1}^{S_{1}}\frac{(i+1/2)}{i^{2}}+S-S_{1}\approx S_{1}\ln(S_{1})+S,$ (34)
for $S>S_{1}$. In the same way, the value of $F_{S}=\sum_{i=1}^{S}\nu_{i}F_{Q_{i}}$ can be approximated by $F_{S}=\sum_{i=1}^{S}\nu_{i}^{\textrm{min}}F_{Q_{i}}$. If one considers the case where each measurement is done on a product state with $N$ qubits, then $F_{Q_{i}}=N\pi^{2}i^{2}/4$. Notice that, in this case, the value of $\nu_{i}^{\textrm{min}}$ is given by the value in Eq. (33) divided by $N$. The same happens to $S_{1}$. Using Eq. (33), the value of $F_{S}$ is given by
$F_{S}$ $=$ $N\frac{\pi^{2}}{4}\left[\frac{S_{1}}{2N}\sum_{i=1}^{S_{1}/N}(2i+1)+\sum_{i=S_{1}/N+1}^{S}i^{2}\right]$ (35)
$=$ $N\frac{\pi^{2}}{24}\left[\frac{S_{1}}{N}\!\left(\frac{S_{1}^{2}}{N^{2}}\!-\!3\frac{S_{1}}{N}\!-\!1\right)\!\!+\!\!S(S\!+\!1)(2S\!+\!1)\right],$
when $S>S_{1}$. Using Eq. (34) and assuming that $S\gg 1$, the above expression can be rewritten as
$F_{S}\approx N\frac{\pi^{2}}{24}\left[\frac{S_{1}}{N}\!\left(\frac{S_{1}^{2}}{N^{2}}\!-\!3\frac{S_{1}}{N}\!-\!1\right)\!\!+\!\!2\left(\nu-\frac{S_{1}}{N}\ln(\frac{S_{1}}{N})\right)^{3}\right],$ (36)
where the value of $S_{1}$ is given by the value that follows from Eq (33), which corresponds to the case of a single probe. Inserting this expression into Eq. (32), one arrives finally at
$\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]\geq\frac{24(C_{l})^{S}}{N\pi^{2}\left[\frac{S_{1}}{N}\!\left(\frac{S_{1}^{2}}{N^{2}}\!-\!3\frac{S_{1}}{N}\!-\!1\right)\!\!+\!\!2\left(\nu-\frac{S_{1}}{N}\ln(\frac{S_{1}}{N})\right)^{3}\right]}+\left(1-C_{l}\right)\left(\sum_{i=1}^{S}C_{l}^{i-1}E_{i}^{2}\right)\,,$ (37)
for $S>S_{1}$ and $\nu\gg 1$. Assuming that the above inequality can be saturated, the error in the estimation of $\tilde{\omega}$ decreases as $1/\nu^{3}$ under the ATFE, as long as the second term is negligible, which, compared with the non-time-adaptive strategy (cf. Eq. (4)), is a huge advantage. This behavior is shown in Fig. 4. As already noted, the second term does not depend on $\nu$ and dominates when $\nu\to\infty$, limiting the minimum error reachable in the ATFE. A rough upper bound on that term can be stablished as
$\left(1-C_{l}\right)\sum_{i=1}^{S}(C_{l})^{i-1}E_{i}^{2}=\left(1-C_{l}\right)\sum_{i=1}^{S}\frac{(C_{l})^{i-1}}{(i+1)^{2}}\leq\left(1-C_{l}\right)\sum_{i=1}^{\infty}\frac{1}{(i+1)^{2}}\approx 0.64(\left(1-C_{l}\right).$ (38)
In section III.3 it will be numerically shown that the inequality in Eq. (37) can be saturated.
We consider now the case where each measurement is done in a GHZ state with $N_{\textrm{GHZ}}$ qubits. In this situation, as discussed before, $\tilde{t}_{1}=1/(4N_{\textrm{GHZ}})$ and $F_{Q_{i}}=N_{\textrm{GHZ}}^{2}\left(2\pi\tilde{t}_{i}\right)^{2}$. Using $\tilde{t}_{i}=i\tilde{t}_{1}$ results in $F_{Q_{i}}=\pi^{2}i^{2}/4$. This is equal to the Fisher information obtained with measurements done on a single probe. Consequently, the lower bound on the error in the estimation of $\tilde{\omega}$ with GHZ states of $N_{\textrm{GHZ}}$ probes is given by substituting $N=1$ in Eq. (37), which is an improvement over Eq. (7). On the other side, this implies that, if only the total number of probes is counted as resource, product states are better than GHZ states even when using time-adaptive methods.
Another resource to be taken into account is the total duration of the experiment $T=\sum_{i=1}^{S}\nu_{i}^{\textrm{min}}\tilde{t}_{i}=\sum_{i=1}^{S_{1}}\nu_{i}^{\textrm{min}}\tilde{t}_{i}+\sum_{i=S_{1}+1}^{S}\tilde{t}_{i}$. For measurements on an $N$-probe GHZ state, $\nu_{i}^{\textrm{min}}$ and $S_{1}$ are given by Eq. (33) and $\tilde{t}_{i}=i/(4N)$, as discussed above. In this case
$T$ $=$ $\frac{S_{1}}{4N}\sum_{i=1}^{S_{1}}\frac{i+1/2}{i}+\frac{1}{4N}\sum_{i=S_{1}+1}^{S}i$ (39)
$\approx$ $\frac{1}{8N}\left[2S_{1}\left(S_{1}+\ln(S_{1})-1\right)+S(S+1)\right].$
When $S\gg 1$, using Eq. (34) and setting $N=1$ in Eq. (37), one can see that the bound on $\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]$ scales as $1/(NT)^{3/2}$, instead of as $1/(NT)$, which is the bound on the error without the use of time-adaptive strategies (see Eqs. (5)).
For measurements on an N-probe product state, $\nu_{i}^{\textrm{min}}$ is equal to the value given by Eq. (33) divided by N, $S_{1}\to S_{1}/N$ and $\tilde{t}_{i}=i/4$, leading to
$T$ $=$ $\frac{S_{1}}{4N}\sum_{i=1}^{S_{1}/N}\frac{i+1/2}{i}+\frac{1}{4}\sum_{i=S_{1}/N+1}^{S}i$ (40)
$\approx$ $\frac{1}{8}\left[2\frac{S_{1}}{N}\left(\frac{S_{1}}{N}+\ln(\frac{S_{1}}{N})-1\right)+S(S+1)\right].$
Now, for $S\gg 1$, the use of Eq. (34) shows the that the bound on $\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]$ scales as $1/(NT^{3/2})$. This scaling with $NT$ is worser than the scaling for GHZ states. Nevertheless, it is worth to recall that, given $T$, the number of measurements when using GHZ states is $\sqrt{N}$ times the number of measurements on product states. This, in turn, means that the total number of probes used in the estimation with GHZ states is $\sqrt{N}$ times the total number of probes used in the estimation with product states. Consequently, if $N^{3/2}$- probe product states are used, instead of N- probe product states, the same variance on the estimation of $\tilde{\omega}$ is reached, for given $T$, as the variance obtained with N-probe GHZ states, using, in both cases, the same total number of probes. Therefore, if the total number of probes and the total time $T$ are counted as resources, there is no advantage in the use of GHZ sates, when compared with the use of product states. This will be shown numerically in section III.3.
In order to obtain a simple analytical expression, the bound in Eq. (32), exemplified in Fig. 3, was derived under the assumption that the confidence levels of the confidence intervals $CI_{i}$, obtained at the primary steps $i$ of the ATFE, were equals and fixed. However, after the step $i=S_{1}$, the Fisher information obtained in a single measurement is larger than the Fisher information necessary to produce a confidence interval with marginal error $E_{i}=1/(i+1)$ and confidence level $C_{l}$. Since the marginal errors $E_{i}$ are fixed, this implies that the confidence level $C_{l}$ increases after the step $i=S_{1}$. This, on the other side, decreases the probability that the parameter is outside the resulting confidence intervals, diminishing the error in the estimation of $\omega$. This was not taken into account in deriving the bound in Eq. (32). Consequently, the actual upper bound on $\text{Var}_{\omega}\left[\widehat{\omega}_{\textrm{MLE}}\right]$ is smaller and tighter than the bound given in that equation. Furthermore, in deriving that bound, the confidence intervals were updated only after $\nu_{i}^{\textrm{min}}$ measurements, inside each primary step $i$, instead of after every measurement. During the numerical simulations, we could see that, by updating the confidence interval after each measurement, using the previous MLE, one obtains smaller errors, since those estimates which were close, but outside the confidence interval, have a chance of being inside the updated interval. Unfortunately, we do not have an analytical expression for the upper bound to the Holevo variance, when the two points above are taken into account. For this reason, we shall rely on numerical simulations for assessing the final error obtainable in the ATFE protocol.
III.3 Numerical simulations of the protocol
We summarize the ATFE strategy for GHZ states in the algorithm 1. For the first series of measurements, correspoding to strategy $i=1$, each with time $\tilde{t}_{1}$, we choose $\nu_{1}$ in AQSE such that the asymptotic behavior of the MLE is reached (i.e. the distribution of the estimator is approximately normal), this allow us to use the formula Eq. (25). From strategies $i>1$, we use the minimum number of measurements, $\nu_{i}^{\textrm{min}}$, given by Eq. (33), as the distribution of the estimator is already approximately normal since strategy $i=1$. For confidence levels close to one (such as $0.999$), we numerically observe that the estimator is approximately normal after $\nu_{1}^{\textrm{min}}$ measurements (see Eq. (33)).
Input:
Confidence: $1-\alpha$,
Number of particles: $N_{\textrm{GHZ}}$,
Number of initial measurements: $\nu_{1}$,
Number of total measurements: $\nu\geq\nu_{1}$.
Output:
Estimate $\widehat{\omega}(x)$ after $\nu$ measurements.
1 Initialize variable: $x\leftarrow x_{0}\in\left\{0,1\right\}$;
2 Initialize parameter space: $\tilde{\Omega}_{0}=[-1,1)$;
3 First guess: $\widehat{\omega}(x)\leftarrow\textrm{rand}(1,[-1,1))$;
4 $i\leftarrow 1$;
5 $\tilde{t}_{1}\leftarrow 1/4N_{\textrm{GHZ}}$;
6 $L(x_{0};\tilde{\omega})\leftarrow 1$;
7 for $j=1\text{ to }\nu$ do
8
9      $x_{j}\leftarrow\textrm{ outcome from }P_{\textrm{L}}\left(\widehat{\omega}(x),\tilde{t}_{j}\right)$;
10       $L(x_{j};\tilde{\omega})\leftarrow\textrm{Tr}\left[P\left(x_{j};\widehat{\omega}(x),\tilde{t}_{j}\right)\rho(\tilde{\omega})\right]$;
11       $x\leftarrow x||x_{j}$;
12       $L(x;\tilde{\omega})\leftarrow\prod_{x_{n}\in x}L(x_{n};\tilde{\omega})$;
13       if  $j\geq\nu_{1}$  then
14             $\widehat{\omega}(x)\leftarrow\arg\max_{\Omega_{j-1}}L(x;\tilde{\omega})$;
15
16      else
17             $\widehat{\omega}(x)\leftarrow\arg\max_{\Omega_{0}}L(x;\tilde{\omega})$;
18
19       end if
20      $\tilde{\Omega}_{j}\leftarrow CI(\widehat{\omega}(x);\alpha)$;
21       if  $\mathrm{E}_{i}(\tilde{\Omega}_{j})\leq\frac{1}{i+1}$ AND $j\geq\nu_{1}$  then
22             $i\leftarrow i+1$;
23             $\tilde{t}_{j}\leftarrow\tilde{t}_{j-1}+1/4N_{\textrm{GHZ}}$;
24
25      else
26             $\tilde{t}_{j+1}\leftarrow\tilde{t}_{j}$;
27
28       end if
29
30 end for
31return $\widehat{\omega}(x)$
Algorithm 1 Adaptive-Time Frequency Estimation (ATFE)
We summarize the ATFE strategy for product states in the algorithm 2. In order to compare with the GHZ case, when a product state $\rho^{\otimes N}$, consisting of $N$ qubit states is employed as a probe state, we assume that they are measured at the same time (i.e. in parallel). Note that in this case we have to choose the POVM that we will use in each of the $N$ qubits. For the first strategy, $i=1$, we found better results if we use a randomly generated $\tilde{g}$ (from $-1$ to $1$) for the POVM of each qubit, as this minimizes non-identifiability problems. Subsequently, the strategy operates similarly to the GHZ state case, using the same POVM for each qubit.
Input:
Confidence: $1-\alpha$,
Size of product state: $N$,
Number of initial measurements: $\nu_{1}$,
Number of total measurements: $\nu\geq\nu_{1}$.
Output:
Estimate $\widehat{\omega}(x)$ after $\nu$ measurements.
1 Initialize variable: $x\leftarrow x_{0}\in\left\{0,1\right\}$;
2 Initialize parameter space: $\tilde{\Omega}_{0}=[-1,1)$;
3 First parameters for the POVMs: $\tilde{g}\in\mathbb{R}^{N}\leftarrow\textrm{rand}(N,[-1,1))$;
4 $i\leftarrow 1$;
5 $\tilde{t}_{1}\leftarrow 1/4$;
6 $L(x_{0};\tilde{\omega})\leftarrow 1$;
7 for $j=1\text{ to }\nu$ do
8       do in parallel
9             for $1\leq k\leq N$ do
10                   $x_{k}\leftarrow\textrm{ outcome from }P_{\textrm{L}}\left(\tilde{g}[k],\tilde{t}_{j}\right)$;
11                   $L(x_{k};\tilde{\omega})\leftarrow\textrm{Tr}\left[P\left(x_{k};\tilde{g}[k],\tilde{t}_{j}\right)\rho(\tilde{\omega})\right]$;
12
13             end for
14
15       end
16
17      $x\leftarrow(x_{1},x_{2},\ldots,x_{N})\mathbin{\|}x$;
18       $L(x;\tilde{\omega})\leftarrow\prod_{x_{n}\in x}L(x_{n};\tilde{\omega})$;
19       if  $j\geq\nu_{1}$  then
20             $\widehat{\omega}(x)\leftarrow\arg\max_{\tilde{\Omega}_{j-1}}L(x;\tilde{\omega})$;
21
22      else
23             $\widehat{\omega}(x)\leftarrow\arg\max_{\tilde{\Omega}_{0}}L(x;\tilde{\omega})$;
24
25       end if
26      $\tilde{g}\leftarrow\textrm{rep}\left(\widehat{\omega}(x),N\right)$;
27       $\Omega_{j}\leftarrow CI(\widehat{\omega}(x);\alpha)$;
28       if  $\mathrm{E}_{i}(\tilde{\Omega}_{j})\leq\frac{1}{i+1}$ AND $j\geq\nu_{1}$  then
29             $i\leftarrow i+1$;
30             $\tilde{t}_{j+1}\leftarrow\tilde{t}_{j}+1/4$;
31
32      else
33             $\tilde{t}_{j+1}\leftarrow\tilde{t}_{j}$;
34
35       end if
36
37 end for
return $\widehat{\omega}(x)$
Algorithm 2 Adaptive-Time Frequency Estimation (ATFE) in Parallel
We now present the results obtained from the simulation of the algorithms for ATFE. First we focus on the case of $N=1$ qubit, (in this case algorithm 1 or 2 are the same). The error as a function of the number of measurements is shown in Fig. 4. The solid gray line represents the estimation error predicted by the QCRB for $\tilde{t}=1/4$, which corresponds to the case without adaptive measurements; an estimation error below that line shows the advantage of the adaptive method. The dark blue rhombus represent the lower bound for the estimation error of the ATFE strategy without update of the confidence interval center (see Eq. (32)). The green points corresponds to the numerical simulation of that case, which shows that the QCRB for the non-time-adaptive strategy ($\tilde{t}=1/4$) can be improved. As predicted by Eq. (32), for a large number of measurements ($\nu$), the contribution to the error due to the estimations outside the confidence intervals will eventually dominate. The light blue rhombus in Fig. 4 represent the smallest possible error, which is the QCRB ($C_{l}=1$). Our goal is to achieve an estimation error that is as close as possible to this bound. The results obtained when the confidence interval is updated at each step of the simulation are shown by the red points in Fig. 4. The error is found to be close to the smallest possible error for $\nu$ up to $60$, indicating the large advantage of the proposed measurement strategy. For the case of product states, increasing the number of qubits diminishes the error by a factor of $N$, as predicted by the QCRB.
Figure 4: ATFE Holevo variance vs the number of measurements for one qubit and a confidence level of $0.999$. The lower bound error of the AQSE method for $\tilde{t}=1/4$ (gray line), for the ATFE without center update (green points) and for the ideal case of $C_{l}=1$ (light blue rhombus) are shown as reference. The dark blue rhombus is the error of the simulation of the ATFE strategy with no center update of the confidence interval at each step, it can be seen how the method can be used to diminish the estimation error for $\nu<40$. The red points is the error in the estimation of the simulation of the ATFE strategy with center update of the confidence intervals at each step, it can be seen how the method gives an estimation error close to the ideal bound for $\nu$ up to $60$. The expected Holevo variance at each point for both ATFE strategies was obtained from the average of 5 Monte Carlo simulations of the experiment, each with $10^{3}$ samples, except for points $40$, $50$, and $60$ for ATFE w/o update strategy, where $10^{4}$ samples were used.
We now study what happens when GHZ states are used as probes. Fig. 5 shows the error in frequency estimation using the algorithm 1 for ATFE with $N_{\rm{GHZ}}=1,5,10$ and $C_{l}=0.999$. Again, the solid green line represents the estimation error predicted by the QCRB for $\tilde{t}=1/4$, and the dark blue rhombus represent the smallest possible error, which is the QCRB ($C_{l}=1$). The pink points show the error of AQSE for $\tilde{t}=1/4$, the red, light green, and orange points the estimation error for ATFE with $N_{\rm{GHZ}}=1,3,5$ respectively. As we anticipated from Eq. (7) and from Eq. (37) with $N=1$, the estimation error does not improve by increasing the number of qubits in a GHZ state.
Figure 5: Holevo variance vs the number of measurements for different $N_{\rm{GHZ}}$. The error of ATFE, with update of the confidence interval at each step, for different number of particles $N_{\rm{GHZ}}=1,5,10$ and initial time $\tilde{t}_{1}=1/(4N_{\rm{GHZ}})$. Increasing the number of particles in the GHZ state does not improve the error as a function of the number of performed experiments $\nu$. The error of the AQSE method for $\tilde{t}=1/(4N_{\rm{GHZ}})$ (gray line) and for the ideal case of $C_{l}=1$ (dark blue dots) are shown as reference. The expected Holevo variance at each point for each ATFE strategies was obtained from the average of 5 Monte Carlo simulations of the experiment, each with $10^{3}$ samples,
III.4 Product states vs GHZ states
By using the maximum time that ensures an identifiable estimator at any measurement, one obtains larger errors when GHZ states are the probe states instead of product states. This fact can be seen by comparing equation Eq. (4) with equation Eq. (7), where it is clear that, at each measurement, one learns more about the parameter using product states than with GHZ states. This advantage remains even when time-adaptive strategies are employed. The bound given in Eq. (37), obtained without updating the $\widehat{CI}$ at each measurement step, confirms this prediction, since the use of initial GHZ states with arbitrary number of probes corresponds to putting $N=1$ in that bound. Numerical simulations of the frequency estimation, where the confidence interval is updated at subsequent adaptive steps (with the updating process taking place after the $\nu_{1}$ initial steps), further validate that GHZ states do not outperform product states when considering the total number of qubits used in the experiment (see Fig. 5).
Figure 6: Comparison of total number of qubits and variance for GHZ and product states in parallel for ATFE at $0.999\%$ of confidence. The ratios between the total number of particles and the corresponding QCRBs are examined as a function of time. The comparison encompasses the utilization of GHZ states and product states in parallel for ATFE. Each color corresponds to a distinct dimension of the Hilbert space, with red representing 5 dimensions, green representing 10 dimensions, and blue representing 15 dimensions. The solid line illustrates the ratio of the number of particles, while the dashed line depicts the ratio of variances.
On the other hand, when considering the minimization of the estimation error with respect to the total experiment time $T$, the ATFE protocol using GHZ states exhibits smaller estimator variance compared to product states. This behavior is in accordance with the predictions given by Eqs. (37), (39) and (40), in the case where the confidence interval is not updated at each increment of time. When the confidence interval is updated at each adaptive step, the GHZ states give smaller errors than product states, confirming those predictions. This is shown in Fig. 6, where the ratio between the estimator variance when the probe is a $N$-product state and the estimator variance for a GHZ state with $N$ qubits is plotted. This ratio scales as $\sqrt{N}$ when the total time $T$ is large enough, as predict by Eqs. (39) and (40). However, as already discussed, for a fixed time $T$, the total number of qubits employed using GHZ states is larger by $\sqrt{N}$ than the total number of qubits employed with product states, as more measurements are needed with GHZ states. In Fig. 6 we also plot the ratio between the total number of probes used for GHZ states and for product states. Specifically, we explore scenarios involving $N_{\rm{GHZ}}=5,10,15$ for ATFE with GHZ states, as well as a set of $N=5,10,15$ product states $\rho(\omega)^{\otimes N}$, which are measured in parallel at each adaptive step. In every case, we maintain a confidence level of $C_{l}=0.999$ and set $\nu_{1}=20$ to ensure asymptotic normality in the distribution of the MLE after the first strategy ($i=1$). It can be seen that the ratio between the variances tends to the same value as the ratio between the number of particles. This means that $\sqrt{N}$ more qubits were used with GHZ states than with product states. On the other side, since increasing the total number of probes by $\sqrt{N}$, for product states, decreases the estimation variance by the same factor, this implies that, when the total number $N_{T}$ of probes and the total time $T$ of the estimation are fixed, product states and GHZ have the same performance for the time-adaptive strategy presented here, in contrast to the non-adaptive strategy, where the use of product states is more advantageous than the use of GHZ states.
IV Discussion & Conclusions
In this article, we investigated the real advantages of the use of entangled states for improving frequency estimation, with special focus on the spectroscopy of two-level systems. The basic metrological scheme to estimating a transition frequency $\omega$, using $N$ two-level probes, consists in first preparing the $N$ probes in an adequate quantum state, letting them evolve for a time $t$ and then measuring some observable on the final state. The measurement results are fed into an estimator function, which produces an estimate of $\omega$. The QCRB provides the ultimate lower bound on the uncertainty of the estimate of the value of $\omega$, and in principle, this bound can be reached using the maximum likelihood estimator. That bound shows that, when the estimation procedure consists of a large number of preparation-sensing cycles, with a fixed sensing time and a fixed number of probes in each run of the experiment, the use of probes in a maximally entangled state leads to an enhanced precision in the frequency estimation when compared with the use of the same number of probes in a product state. However, in order to that bound be reachable it is necessary that the statistical model be identifiable. This, in turn, puts a limit on the maximal sensing time allowed in each run of the estimation procedure. We have discussed in detail the effects of this restriction on the maximal sensing time on the de facto reachable bound on the precision of the frequency estimation and, in particular, how the requirement on the identifiability produces a maximum likelihood estimator that annuls any possible advantage of the use of maximally entangled probes when compared with use of independent probes. In fact, when the total number of probes and the total sensing time are counted as resources, it is more advantageous to use independent probes than maximally entangled ones.
As a means to counteract the limitations on the maximal sensing time in frequency estimation, imposed by the requirement of statistical model identifiability, we presented a time-adaptive estimation strategy. In this strategy, one assumes that some prior knowledge about the value of the frequency $\omega$ already exists and that this value lies inside some frequency interval of a given length. Using the maximal sensing time allowed by that interval length, one starts with a small number of identical preparation-sensing-measurement cycles to produce a first estimation of $\omega$. This first estimation allows one to shorten the interval length and, consequently, to increase the sensing time for the next estimation cycle. This process is repeated until the desired estimation uncertainty is reached. We have determined a reachable error bound for the presented strategy and discussed how to optimally choose its parameters in order to minimize that bound. The bound shows that the time-adaptive strategy leads to much better scaling of the estimation uncertainty with the total number of probes and the total sensing time than the traditional fixed-sensing-time strategy. The bound also shows that, when the total number of probes and the total sensing time are counted as resources, independent probes and maximally entangled ones have now the same performance, in contrast to the non-adaptive strategy, where the use of product states is more advantageous than the use of maximally entangled states. Finally, we presented numerical simulations of the proposed time-adaptive estimation strategy. Those simulations confirmed all the analytical predictions presented in this article.
Acknowledgments
We thank Laboratorio Universitario de Cómputo de Alto Rendimiento (LUCAR) of IIMAS-UNAM for their service on information processing. MR acknowledges funding from the National Science Foundation (NSF) grants FRHTP # PHY-2116246, and # 2016244 - QLCI-CI. PB-B gratefully acknowledges support from the PAPIIT-DGAPA grant IG101324.